OpenStudy (anonymous):

Factor out the GCF, if possible -6x^5y^5 + 30x^4y^4 - 12x^3y^3

OpenStudy (mathteacher1729):

Let's take this apart bit by bit. First, the integers: What is the GCF of (6, 30, and 12) ? What is the GCF of (x^5 , x^4 and x^3)? What is the GCF of (y^5, y^4 and y^3) ? Once you have that, then you're ready for the next step. :)

OpenStudy (anonymous):

I dont get it...

OpenStudy (anonymous):

:( I'll just try to do it myself if no one can help..

OpenStudy (mathteacher1729):

Do you understand what GCF means? That is critical to this problem.

OpenStudy (turingtest):

What is the GCF of (x^5 , x^4 and x^3) is a good example. The GFC of these is x^3 because this is a factor of all the others: x^5=x^3(x^2) x^4=x^3(x) x^3=x^3(1) or use 12 and 16: 12=4*3 16=4*4 they both share a 4, so the GFC of 12 and 16 is 4

OpenStudy (anonymous):

\[-6 x^5 y^5+30 x^4 y^4-12 x^3 y^3=-6 x^3 y^3 \left(x^2 y^2-5 x y+2\right) \]